OpenStudy (anonymous):

What is the simplified version of this equation. y=4(x-2)^2+1

jhonyy9 (jhonyy9):

\[y=4(x ^{2}-4x+4)+1=4x ^{2}-16x+16+1=4x ^{2}-16x+17\]

OpenStudy (anonymous):

4x^2 - (16x+4) + 1 =4X^2 - 16x +16+1=4x^2 -16x+17 = 48

jhonyy9 (jhonyy9):

princessgirl16 how you have got it this result ?